# AnyAttack: Targeted Adversarial Attacks on Vision-Language Models Toward Any Images

### v2

@@ Abstract
Due to their multimodal capabilities, Vision-Language Models (VLMs) have found numerous impactful applications in real-world scenarios. However, recent studies have revealed that VLMs are vulnerable to image-based adversarial attacks, particularly targeted adversarial images that manipulate the model to generate harmful content specified by the adversary. Current attack methods rely on predefined target labels to create targeted adversarial attacks, which limits their scalability and applicability for large-scale robustness evaluations. In this paper, we propose AnyAttack, a self-supervised framework that generates targeted adversarial images for VLMs without label supervision, allowing any image to serve as a target for the attack. Our framework employs the “pre-training and fine-tuning” paradigm, with the adversarial noise generator pre-trained on the large-scale LAION-400M dataset. This large-scale pre-training endows our method with powerful transferability across a wide range of VLMs. Extensive experiments on five mainstream open-source VLMs (CLIP, BLIP, BLIP2, InstructBLIP, and MiniGPT-4) across three multimodal tasks (image-text retrieval, multimodal classification, and image captioning) demonstrate the effectiveness of our attack. Additionally, we successfully transfer AnyAttack to multiple commercial VLMs, including Google Gemini, Claude Sonnet, Microsoft Copilot and OpenAI GPT. These results reveal an unprecedented risk to VLMs, highlighting the need for effective countermeasures.
@@ 1 Introduction
Vision-Language Models (VLMs) have exhibited remarkable performance across a diverse array of tasks, primarily attributed to the scale of training data and model size [18, 11, 37]. Despite their remarkable performance, these models, heavily reliant on visual inputs, remain vulnerable to image-based adversarial attacks [^1] For simplicity, we will refer to image-based adversarial attacks as “adversarial attacks” in the remainder of this paper, distinguishing them from text-based adversarial attacks., which are carefully crafted input images designed to mislead the model into making incorrect predictions [22]. While general, untargeted adversarial attacks only aim to induce incorrect outputs, targeted adversarial attacks present a more insidious threat, manipulating the model’s output to yield an adversary-specified, predetermined response. For instance, a benign image such as a landscape could be subtly altered to elicit harmful text descriptions such as “violence” or “explicit content” from the model. Such manipulation could have severe implications for content moderation systems, potentially leading to the removal of legitimate content or the inappropriate distribution of harmful material.
@@
Figure 1: Comparison of existing targeted adversarial attack strategies and the self-supervised adversarial noise pre-training method AnyAttack with fine-tuning capability.
Both targeted adversarial attacks and jailbreak techniques [38, 1, 2] aim to induce harmful responses from models. However, the key distinction lies in their outputs: targeted adversarial attacks produce adversary-specified, predetermined responses, whereas jailbreak attacks elicit non-predetermined responses. As VLMs become increasingly accessible to the public, facilitating the rapid proliferation of downstream applications, this vulnerability poses a significant threat to the reliability and security of VLMs in real-world scenarios. Therefore, exploring new targeted attack methods tailored to VLMs is crucial to address these vulnerabilities. However, existing targeted attack methods on VLMs present challenges due to the reliance on target labels for supervision, which limits the scalability of the training process. For example, it is impractical to expect a generator trained on ImageNet [19] to produce effective adversarial noise for VLMs.
@@
To overcome this limitation, we propose a novel self-supervised framework, AnyAttack, which leverages the original image itself as supervision, enabling any image to serve as a target for the deployment of targeted adversarial attacks. Our approach involves pre-training a generator on the large-scale LAION-400M dataset [20], enabling the pre-trained noise generator to learn comprehensive noise patterns from diverse image data. Through self-supervised adversarial noise pre-training, we can further fine-tune the pre-trained generator on downstream datasets for adapting downstream vision-language tasks. The large-scale pre-training establishes a foundation model (we release the model weights), thereby enhancing the potential for developing more powerful adversarial attacks. Our framework, to the best of our knowledge, is the first to implement the “pre-training and fine-tuning” paradigm for targeted adversarial attacks at scale. Fig. 1 highlights the distinctions between our method and existing strategies.
@@
  • We propose AnyAttack, a self-supervised framework that utilizes the original image as supervision, allowing any image to serve as a target for deploying targeted adversarial attacks.
@@ 2 Related Work
VLMs have revolutionized DNNs by using large-scale pre-training on diverse image-text datasets. Broadly, VLMs can be categorized into two types: the first offers multimodal functionalities built on large language models (LLMs), complemented by visual models, such as BLIP2 [11], InstructBLIP [3], and MiniGPT-4 [37]. The second type provides a more balanced approach, bridging textual and visual modalities efficiently, as seen in models like CLIP [18], ALIGN [9], and BLIP [10]. Recent advancements in VLMs have spurred parallel research into their vulnerabilities, with jailbreaks and adversarial attacks emerging as distinct threat vectors. Multimodal jailbreaks primarily exploit cross-modal interaction vulnerabilities in VLMs, with the intention of influencing LLMs [38]. These attacks manipulate inputs of text [28], images [2, 8, 17, 16], or both simultaneously [32, 23], aiming to elicit harmful but non-predefined responses. In contrast, image-based adversarial attacks focus on manipulating the image encoder of VLMs, typically leaving the LLM component largely undisturbed. The objective is to induce adversary-specified, predetermined responses through precise visual manipulations. Understanding these differences is crucial for explaining our method.
@@ 3.1 Preliminaries
We begin by formulating the problem of targeted adversarial attacks. Let $f_{s}$ represent a pre-trained surrogate model, and $\mathcal{D}=\{(x,y)\}$ denote the image dataset, where $x$ is the original image and $y$ is the corresponding label (description). The attacker’s objective is to craft an adversarial example $x^{\prime}=x+\delta$ that misleads the target model $f_{t}$ into predicting a predefined target label $y_{t}$. In the context of VLMs, this objective requires that $x^{\prime}$ aligns with $y_{t}$ as a valid image-text pair. The process of generating targeted adversarial images typically involves finding a perturbation $\delta$ using the surrogate model $f_{s}$.
@@ 3.2 AnyAttack
To increase the number of random images within every batch, we present the $K$-augmentation strategy, which duplicates both adversarial noises $\delta$ and the original images $x$ $K$ times, forming $K$ mini-batches. For each mini-batch, the order of the adversarial noises remain consistent, while the order of the original images is shuffled within the mini-batch, referred to as shuffled images. These shuffled images are then added to the corresponding adversarial noise, resulting in adversarial images $x^{\prime}$. Next, the adversarial images are fed into $F$ to produce adversarial embeddings $\mathbf{z}^{(adv)}$, which are then used for subsequent calculations against the original embeddings $\mathbf{z}$.
@@
Depending on the downstream tasks, self-supervised adversarial noise fine-tuning employ two different fine-tuning objectives. The first strategy is tailored for the image-text retrieval task, which imposes stricter requirements for distinguishing between similar samples. It demands robust retrieval performance in bi-directions: from $\mathbf{z}^{(adv)}$ to $\mathbf{z}$ and from $\mathbf{z}$ to $\mathbf{z}^{(adv)}$, denoted as $\mathcal{L}_{\text{Bi}}$.
@@ 4.4 Evaluation on Image Captioning
Here, we evaluate the performance of our attack on the image captioning task using the MSCOCO dataset. The VLMs take adversarial images as input and generate text descriptions, which are then assessed against the ground-truth captions using standard metrics. Tab. 3 presents the results across four VLMs: InstructBLIP, BLIP2, BLIP, and MiniGPT-4. Our proposed attack method, AnyAttack-Cos w/ Auxiliary consistently demonstrates superior performance across all evaluation metrics, outperforming the baseline attacks on each VLM.
@@ 4.5 Transfer to Commercial VLMs
We selected 100 images from the MSCOCO dataset as target images and conducted a comparison between our method and baseline approaches. Using the prompt “Evaluate the relationship between the given image and text”, along with corresponding options, Tab. 4 reports the percentage of responses labeled as “relevant” by these commercial VLMs. Our method consistently outperforms the baselines, achieving substantial improvements across all evaluated models.

### v3

# AnyAttack: Towards Large-scale Self-supervised Adversarial Attacks on Vision-language Models

@@ Abstract
Due to their multimodal capabilities, Vision-Language Models (VLMs) have found numerous impactful applications in real-world scenarios. However, recent studies have revealed that VLMs are vulnerable to image-based adversarial attacks. Traditional targeted adversarial attacks require specific targets and labels, limiting their real-world impact. We present AnyAttack, a self-supervised framework that transcends the limitations of conventional attacks through a novel foundation model approach. By pre-training on the massive LAION-400M dataset without label supervision, AnyAttack achieves unprecedented flexibility - enabling any image to be transformed into an attack vector targeting any desired output across different VLMs. This approach fundamentally changes the threat landscape, making adversarial capabilities accessible at an unprecedented scale. Our extensive validation across five open-source VLMs (CLIP, BLIP, BLIP2, InstructBLIP, and MiniGPT-4) demonstrates AnyAttack’s effectiveness across diverse multimodal tasks. Most concerning, AnyAttack seamlessly transfers to commercial systems including Google Gemini, Claude Sonnet, Microsoft Copilot and OpenAI GPT, revealing a systemic vulnerability requiring immediate attention.
@@ 1 Introduction
Figure 1: Comparison of existing targeted adversarial attack strategies and the our proposed self-supervised method - AnyAttack.
Vision-Language Models (VLMs) have exhibited remarkable performance across a diverse array of tasks, primarily attributed to the scale of training data and model size [16, 9, 35]. Despite their remarkable performance, these models, heavily reliant on visual inputs, remain vulnerable to image-based adversarial attacks [^1] For simplicity, we will refer to image-based adversarial attacks as “adversarial attacks” in the remainder of this paper, distinguishing them from text-based adversarial attacks., which are carefully crafted input images designed to mislead the model into making incorrect predictions [20]. The evolution of adversarial attacks has progressed from general untargeted attacks (causing arbitrary errors) to more concerning targeted attacks, where adversaries can manipulate VLMs to produce specific, predetermined harmful outputs. For instance, a benign image such as a landscape could be subtly altered to elicit harmful text descriptions such as “violence” or “explicit content” from the model. Such manipulation could have severe implications for content moderation systems, potentially leading to the removal of legitimate content or the inappropriate distribution of harmful material.
@@
As VLMs become increasingly accessible to the public, facilitating the rapid proliferation of downstream applications, this vulnerability poses a significant threat to the reliability and security of VLMs in real-world scenarios. Therefore, exploring new targeted attack methods tailored to VLMs is crucial to address these vulnerabilities. However, existing targeted attack methods on VLMs present challenges due to the reliance on target labels for supervision, which limits the scalability of the training process. For example, it is impractical to expect a generator trained on ImageNet [17] to produce effective adversarial noise for VLMs. To overcome this limitation, we propose a novel self-supervised framework, AnyAttack, which leverages the original image itself as supervision, enabling any image to be transformed into an attack vector targeting any desired output across different VLMs. Our approach involves pre-training a generator on the large-scale LAION-400M dataset [18], enabling the pre-trained noise generator to learn comprehensive noise patterns from diverse image data. Through self-supervised adversarial noise pre-training, we can further fine-tune the pre-trained generator on downstream datasets for adapting downstream vision-language tasks. The large-scale pre-training establishes a foundation model, thereby enhancing the potential for developing more powerful adversarial attacks. Our framework, to the best of our knowledge, is the first to implement the “pre-training and fine-tuning” paradigm for targeted adversarial attacks at scale, breaking the barriers of traditional attack methods. Fig. 1 highlights the distinctions between our method and existing strategies.
@@
  • We propose AnyAttack, a self-supervised framework that utilizes the original image as supervision, allowing any image to be transformed into an attack vector targeting any desired output across different VLMs.
@@ 2 Related Work
Multimodal jailbreaks primarily exploit cross-modal interaction vulnerabilities in VLMs. These attacks manipulate inputs of text [26], images [1, 7, 15, 14], or both simultaneously [30, 21], aiming to elicit harmful but non-predefined responses. In contrast, image-based adversarial attacks focus on manipulating the image encoder of VLMs. The objective is to induce adversary-specified, predetermined responses through precise visual manipulations.
@@ 3.1 Preliminaries and Adversary’s Settings
We begin by formulating the problem of targeted adversarial attacks. Let $f_{s}$ represent a pre-trained surrogate model, and $\mathcal{D}=\{(x,y)\}$ denote the image dataset, where $x$ is the original image and $y$ is the corresponding label (description). The attacker’s objective is to craft an adversarial example $x^{\prime}=x+\delta$ that misleads the target model $f_{t}$ into predicting a predefined target label $y_{t}$. In the context of VLMs, this objective requires that $x^{\prime}$ aligns with $y_{t}$ as a valid image-text pair. The process of generating targeted adversarial images typically involves finding a perturbation $\delta$ using the surrogate model $f_{s}$. Tab. 1 highlights the key differences between our approach and current strategies, where $x_{r}$ denotes an irrelevant random image.
Strategy Formulation Target label supervision $\min\mathcal{L}(f_{s}(x+\delta),y_{t}),\quad\text{s.t.}\ y_{t}\neq y$ Target image supervision $\min\mathcal{L}(f_{s}(x+\delta),f_{s}(x_{t})),\quad\text{s.t.}\ x_{t}\neq x$ AnyAttack (ours) $\min\mathcal{L}(f_{s}(\delta+x_{r}),f_{s}(x))$
Table 1: The formulation of different attack strategies. The existing strategies rely on explicit target supervision, whereas our AnyAttack is unsupervised.
@@ 3.2 AnyAttack
To increase the number of random images within every batch, we present the $K$-augmentation strategy, which duplicates both adversarial noises $\delta$ and the original images $x$ $K$ times, forming $K$ mini-batches. For each mini-batch, the order of the adversarial noises remains consistent, while the order of the original images is shuffled within the mini-batch, referred to as shuffled images. These shuffled images are then added to the corresponding adversarial noise, resulting in adversarial images $x^{\prime}$. Next, the adversarial images are fed into $F$ to produce adversarial embeddings $\mathbf{z}^{(adv)}$, which are then used for subsequent calculations against the original embeddings $\mathbf{z}$.
@@
Depending on the downstream tasks, self-supervised adversarial noise fine-tuning employs two different fine-tuning objectives. The first strategy is tailored for the image-text retrieval task, which imposes stricter requirements for distinguishing between similar samples. It demands robust retrieval performance in bi-directions: from $\mathbf{z}^{(adv)}$ to $\mathbf{z}$ and from $\mathbf{z}$ to $\mathbf{z}^{(adv)}$, denoted as $\mathcal{L}_{\text{Bi}}$.
@@ 4.4 Evaluation on Image Captioning
Here, we evaluate the performance of our attack on the image captioning task using the MSCOCO dataset. The VLMs take adversarial images as input and generate text descriptions, which are then assessed against the ground-truth captions using standard metrics. Tab. 4 presents the results across four VLMs: InstructBLIP, BLIP2, BLIP, and MiniGPT-4. Our proposed attack method, AnyAttack-Cos w/ Auxiliary, consistently demonstrates superior performance across all evaluation metrics, outperforming the baseline attacks on each VLM.
@@ 4.5 Transfer to Commercial VLMs
We selected 100 images from the MSCOCO dataset as target images and conducted a comparison between our method and baseline approaches. The experiments were conducted using Google Gemini (Gemini 1.5 Flash) and OpenAI GPT (GPT-4o mini) via their respective APIs. Using the prompt “Evaluate the relationship between the given image and text”, along with corresponding options “A) The text is highly relevant to the image. B) The text is partially relevant to the image. C) The text is not relevant to the image.” Tab. 5 reports the percentage of responses labeled as “highly and partially relevant” by these commercial VLMs. Our method consistently outperforms the baselines, achieving substantial improvements across all evaluated models.
@@ 5 Conclusion
6 Acknowledgements
This work is supported by the National Key R&D Program of China (Grant No. 2023YFC3310700, 2022ZD0160103) and the National Natural Science Foundation of China (Grant No. 62276067). This research has been made possible by a Research Impact Fund project (RIF R6003-21) and a General Research Fund project (GRF 16203224) funded by the Research Grants Council (RGC) of the Hong Kong Government.